I need help verifying trid identities. The problem is (1 - sinx)/(1 + sinx) = (secx - tanx) squared. I multiplied by 1 - sin and got cos squared on bottom, then i got stuck.
Answer by Alan3354(69443)

(1 - sinx)/(1 + sinx) = (secx - tanx)^2
(1 - sinx)/(1 + sinx) = (1/cos - sin/cos)^2
(1 - sinx)/(1 + sinx) = ((1-sin)/cos)^2
(1 - sinx)^2/(1 - sin^2x) = ((1-sin)/cos)^2
(1-sin)^2/cos^2 = (1-sin)^2/cos^2 They're equal